You have $10,000 to invest in a stock portfolio. Your choices are Stock X with an...

  1. You have $10,000 to invest in a stock portfolio. Your choices are Stock X with an expected return of 13 percent and Stock Y with an expected return of 8.6 percent. If your goal is to create a portfolio with an expected return of 11.9 percent, how much money will you invest in Stock X?

    2500

    5000

    7500

    7600
Answer #1

Let weight of Stock X be x and weight of Stock Y be (1 - x)

Expected Return of Portfolio = Weight of Stock X * Expected Return of Stock X + Weight of Stock Y * Expected Return of Stock Y
0.1190 = x * 0.1300 + (1 - x) * 0.0860
0.1190 = x * 0.1300 + 0.0860 - x * 0.0860
0.0330 = x * 0.0440
x = 0.75

Weight of Stock X = 0.75
Weight of Stock Y = 1 - 0.75
Weight of Stock Y = 0.25

Investment in Stock X = 0.75 * $10,000
Investment in Stock X = $7,500